# Using LLMs as Speech-to-Text Retrieval Systems

### v2

@@ 1 Introduction
We posit that we can retrieve speech and text utterances by aligning both modalities within the embedding space built from a single pre-trained LLM. We take inspiration from previous works that use pre-trained LLMs to perform automatic speech recognition (ASR) and automatic speech translation (AST) Rubenstein et al. (2023); Wang et al. (2023); Hassid et al. (2023). Our intuition is that we can perform the speech and text alignment leveraging the capabilities of text-only LLMs without requiring two separate models.
@@
    We show that our model exhibits cross-lingual speech and text matching without training on this type of data. At the same time, we find that cross-lingual speech and text matching is furthered improved by training on readily available machine translation data.
@@ 2.1 Generating Audio Tokens
We convert raw speech into discrete tokens using the process in Lakhotia et al. (2021); Borsos et al. (2023). The process converts a speech query $x_{i}$ into an embedding using a pre-trained speech encoder. The output embedding is then discretized into a set of tokens using k-means clustering. We refer to the resulting tokens as audio tokens. We use the 2B variant of the Universal Speech Model (USM) encoder Zhang et al. (2023) as the speech encoder and take the middle layer as the embedding for $x_{i}$. Additionally, we generate audio tokens at 25Hz using k-means clustering, resulting in a set of 1024 possible audio tokens. We will refer to this as our audio token vocabulary.
@@ 5 Experiments
PaLM 2 DE (Proposed Model) 86.15 13.85
@@ 5.1.1 Seen-Unseen Breakdown
In Figure 2 we break down the $R@1$ scores based on seen and unseen languages during training. We find that our model performs best on the 20 languages that are within the training and evaluation data, but still perform significantly well on the remaining 82 unseen languages. We hypothesize this is due to the vast textual multilingual data our backbone LLM has seen during pre-training.
@@ 5.1.2 Language Group Breakdown
Language Group (#) mSLAM DE PaLM 2 DE
@@
Afro-Asiatic (7) 73.67 90.82
Atlantic-Congo (15) 86.77 79.47
Austro-Asiatic (3) 47.90 41.71
Austronesian (6) 75.50 85.74
Dravidian (4) 65.70 90.46
Indo-European (51) 84.62 92.38
Japonic (1) 5.80 63.23
Kartvelian (1) 70.50 74.57
Koreanic (1) 5.20 45.81
Kra-Dai (1) 3.20 37.12
Mongolic (1) 70.70 98.10
Nilo-Saharan (1) 91.00 94.53
Sino-Tibetan (3) 3.40 61.91
Turkic (4) 81.28 91.89
Uralic (3) 91.40 93.38
Table 2: FLEURS S2T (R@1) performance broken down by language groups. Bold represents better performance. Numbers in parenthesis represent the number of languages within the language group.
@@ 5.2 Evaluating on Cross-Modal and Cross-Lingual Tasks
We evaluate on S2TT to gauge the cross-modal and cross-lingual capabilities of our model. We show that we can further improve S2TT by simply training on a mixture of S2T and translation data without using any S2TT training data.
@@ 6 Related Work
On the other hand, several works aim to build joint speech and text representations. Chung et al. (2021) introduced w2v-bert which combines masked language modeling and contrastive learning to create speech representations. Bapna et al. (2022) jointly pre-trains on speech and text from unsupervised speech and text data. Recently, Duquenne et al. (2023) employed separate speech and text encoders to generate embeddings in over 200 languages. Nevertheless, there is still a lack of understanding of whether joint speech and text representations can be built from a single encoder. We fill this gap by using pre-trained LLMs to jointly train on speech samples and their transcriptions to show that our approach is capable of speech-text matching in 102 languages.
@@ 8 Acknowledgements
We would like to thank Shankar Kumar and Ankur Bapna for the valuable feedback on the draft of the paper. Chris Tar, Mario Guajardo-Céspedes, and Jason Riesa for the early experiment discussions and feedback. Christian Frank, Duc Dung Nguyen, Alex Tudor, and Dalia El Badawy for helping answer questions about AudioPaLM.

### v3

@@ 1 Introduction
We posit that we can retrieve speech and text utterances by aligning both modalities within the embedding space built from a single pre-trained LLM. We take inspiration from previous works that use pre-trained LLMs to perform automatic speech recognition (ASR) and automatic speech translation (AST) Rubenstein et al. (2023); Wang et al. (2023); Hassid et al. (2023); Gong et al. (2023); Peng et al. (2023). Our intuition is that we can perform the speech and text alignment leveraging the capabilities of text-only LLMs without requiring two separate models.
@@
    We show that our model exhibits cross-lingual speech and text matching without training on this type of data. At the same time, we find that cross-lingual speech and text matching is further improved by training on readily available machine translation data.
@@ 2.1 Generating Audio Tokens
We convert raw speech into discrete tokens using the process in Lakhotia et al. (2021); Borsos et al. (2023). The process converts a speech query $x_{i}$ into an embedding using a pre-trained speech encoder. The output embedding is then discretized into a set of tokens using k-means clustering. We refer to the resulting tokens as audio tokens. We use the 2B variant of the Universal Speech Model (USM) encoder Zhang et al. (2023) as the speech encoder and take the middle layer as the embedding for $x_{i}$. Additionally, we generate audio tokens at 25Hz using k-means clustering  [^1] We use the USM-v2 audio tokenizer from Rubenstein et al. (2023). We will refer to this as our audio token vocabulary.
@@ 5 Experiments
PaLM 2 DE (Proposed Model) 86.7 13.4
@@ 5.1.1 Seen-Unseen Breakdown
In Figure 2 we break down the R@1 scores based on seen and unseen languages during training. We find that our model performs best on the 20 languages that are within the training and evaluation data, but still perform well on the remaining 82 unseen languages. We hypothesize this is due to the vast textual multilingual data our backbone LLM has seen during pre-training.
@@ 5.1.2 Language Group Breakdown
Language Group (#) mSLAM DE PaLM 2 DE ##\## Wins
@@
Afro-Asiatic (7) 73.67 84.22 5
Atlantic-Congo (14) 86.77 70.41 1
Austro-Asiatic (2) 47.90 34.42 0
Austronesian (6) 75.50 90.73 6
Dravidian (4) 65.70 92.06 4
Indo-European (51) 84.62 95.32 49
Japonic (1) 5.80 91.54 1
Kartvelian (1) 70.50 82.92 1
Koreanic (1) 5.20 52.36 1
Kra-Dai (2) 3.20 22.09 1
Mongolic (1) 70.70 99.89 1
Nilo-Saharan (1) 91.00 92.52 1
Sino-Tibetan (3) 3.40 90.66 3
Turkic (5) 81.28 92.86 4
Uralic (3) 91.40 99.04 3
All (102) 76.90 86.72 81
Table 2: FLEURS S2T (R@1) performance by language groups. Bold represents better performance. Numbers in parenthesis are the number of languages within the language group. ##\## Wins is the number of languages where PaLM 2 DE outperforms mSLAM in the language group.
@@ 5.2 Evaluating on Cross-Modal and Cross-Lingual Tasks
We evaluate on S2TT to gauge the cross-modal and cross-lingual capabilities of our model. We show we can improve S2TT by simply combining S2T and translation data without S2TT training data.
@@ 6 Related Work
On the other hand, several works aim to build joint speech and text representations Khurana et al. (2022); Gow-Smith et al. (2023). Chung et al. (2021) introduced w2v-bert which combines masked language modeling and contrastive learning to create speech representations. Bapna et al. (2022) jointly pre-trains on speech and text from unsupervised speech and text data. Recently, Duquenne et al. (2023) employed separate speech and text encoders to generate embeddings in over 200 languages. Nevertheless, there is still a lack of understanding of whether joint speech and text representations can be built from a single encoder. We fill this gap by using pre-trained LLMs to jointly train on speech samples and their transcriptions to show that our approach is capable of speech-text matching in 102 languages.
@@ 8 Acknowledgements
We would like to thank Shankar Kumar, Ankur Bapna, and the anonymous reviewers for the valuable feedback on the draft of the paper. Chris Tar, Mario Guajardo-Céspedes, and Jason Riesa for the early experiment discussions and feedback. Christian Frank, Duc Dung Nguyen, Alex Tudor, and Dalia El Badawy for helping answer questions about AudioPaLM.
@@ A.3 Data
A.4 Performance Breakdown By Language
Table 6 includes the PaLM 2 DE R@1 for each language found in FLEURS. We also include the language group from Table 2 and the number of examples found within each S2T test set.
Idx Language Name Code Family ##\## Examples R@1
mSLAM PaLM 2 DE
1 Afrikaans af Indo-European 414 90.1 99.3
2 Amharic am Afro-Asiatic 516 34.1 69.6
3 Arabic ar Afro-Asiatic 427 82.7 98.8
4 Armenian hy Indo-European 929 50.3 89.7
5 Assamese as Indo-European 980 81.5 87.4
6 Asturian ast Indo-European 946 90.1 100.0
7 Azerbaijani az Turkic 918 83.0 98.4
8 Belarusian be Indo-European 955 90.2 97.2
9 Bengali bn Indo-European 911 83.5 84.6
10 Bosnian bs Indo-European 923 95.5 99.8
11 Bulgarian bg Indo-European 657 95.1 100.0
12 Burmese my Sino-Tibetan 870 2.4 19.3
13 Cantonese yue Sino-Tibetan 819 2.4 83.6
14 Catalan ca Indo-European 938 93.2 100.0
15 Cebuano ceb Austronesian 532 79.8 94.9
16 Croatian hr Indo-European 914 98.0 99.8
17 Czech cs Indo-European 720 98.1 99.6
18 Danish da Indo-European 929 94.1 99.9
19 Dutch nl Indo-European 364 95.3 100.0
20 English en Indo-European 647 96.0 99.1
21 Estonian et Uralic 892 95.6 99.9
22 Filipino fil Austronesian 928 73.1 89.1
23 Finnish fi Uralic 916 93.0 98.9
24 French fr Indo-European 675 90.7 100.0
25 Fula ff Atlantic-Congo 649 81.4 81.7
26 Galician gl Indo-European 927 90.9 100.0
27 Ganda lg Atlantic-Congo 705 90.7 75.7
28 Georgian ka Kartvelian 978 70.5 82.9
29 German de Indo-European 841 91.2 100.0
30 Greek el Indo-European 649 81.2 73.2
31 Gujarati gu Indo-European 1000 77.0 95.9
32 Hausa ha Afro-Asiatic 557 84.5 83.1
33 Hebrew he Afro-Asiatic 792 64.0 76.0
34 Hindi hi Indo-European 417 78.0 83.7
35 Hungarian hu Uralic 902 85.3 98.3
36 Icelandic is Indo-European 46 71.7 97.8
37 Igbo ig Atlantic-Congo 869 85.8 64.9
38 Indonesian id Austronesian 684 79.6 99.4
39 Irish ga Indo-European 829 55.1 69.5
40 Italian it Indo-European 857 93.5 100.0
41 Japanese ja Japonic 650 5.8 91.5
42 Javanese jv Austronesian 722 78.0 97.0
43 Kabuverdianu kea Indo-European 859 95.4 99.9
Idx Language Name Code Family ##\## Examples R@1
mSLAM PaLM 2 DE
44 Kamba kam Atlantic-Congo 798 89.7 81.5
45 Kannada kn Dravidian 831 69.0 88.8
46 Kazakh kk Turkic 841 88.7 83.1
47 Khmer km Austro-Asiatic 765 42.1 20.3
48 Korean ko Koreanic 382 5.2 52.4
49 Kyrgyz ky Turkic 974 84.3 88.6
50 Lao lo Kra-Dai 399 37.0 23.3
51 Latvian lv Indo-European 848 97.4 100.0
52 Lingala ln Atlantic-Congo 440 91.2 76.4
53 Lithuanian lt Indo-European 985 96.8 98.2
54 Luo luo Nilo-Saharan 254 91.0 92.5
55 Luxembourgish lb Indo-European 929 80.5 74.6
56 Macedonian mk Indo-European 967 96.1 98.8
57 Malay ms Austronesian 749 77.7 98.7
58 Malayalam ml Dravidian 944 62.3 88.3
59 Maltese mt Afro-Asiatic 918 92.7 76.0
60 Mandarin cmn Sino-Tibetan 944 5.4 100.0
61 Maori mi Austronesian 890 64.7 65.3
62 Marathi mr Indo-European 1005 69.8 82.4
63 Mongolian mn Mongolic 949 70.7 99.9
64 Nepali ne Indo-European 724 66.1 89.6
65 Northern-Sotho nso Atlantic-Congo 738 80.8 70.3
66 Norwegian nb Indo-European 357 91.9 100.0
67 Nyanja ny Atlantic-Congo 745 85.5 63.6
68 Occitan oc Indo-European 968 77.4 99.4
69 Oriya or Indo-European 875 15.7 95.1
70 Oromo om Afro-Asiatic 41 92.7 100.0
71 Pashto ps Indo-European 510 84.8 91.0
72 Persian fa Indo-European 858 85.4 100.0
73 Polish pl Indo-European 758 95.8 99.3
74 Portuguese pt Indo-European 914 91.9 99.9
75 Punjabi pa Indo-European 574 70.6 96.7
76 Romanian ro Indo-European 882 92.0 100.0
77 Russian ru Indo-European 774 93.2 100.0
78 Serbian sr Indo-European 700 97.7 99.1
79 Shona sn Atlantic-Congo 920 84.1 53.9
80 Sindhi sd Indo-European 977 71.8 85.4
81 Slovak sk Indo-European 791 97.6 99.5
82 Slovenian sl Indo-European 834 97.4 100.0
83 Somali so Afro-Asiatic 1007 68.7 86.0
84 Sorani-Kurdish ckb Indo-European 918 80.8 96.7
85 Spanish es Indo-European 907 69.6 100.0
86 Swahili sw Atlantic-Congo 487 91.2 86.2
Idx Language Name Code Family ##\## Examples R@1
mSLAM PaLM 2 DE
87 Swedish sv Indo-European 758 94.2 100.0
88 Tajik tg Indo-European 590 76.3 92.7
89 Tamil ta Dravidian 582 58.0 98.1
90 Telugu te Dravidian 471 73.5 93.0
91 Thai th Kra-Dai 1011 3.2 20.9
92 Turkish tr Turkic 742 84.5 100.0
93 Ukrainian uk Indo-European 750 93.5 99.3
94 Umbundu umb Atlantic-Congo 264 77.3 62.1
95 Urdu ur Indo-European 299 70.6 91.3
96 Uzbek uz Turkic 861 67.6 94.2
97 Vietnamese vi Austro-Asiatic 850 64.5 48.6
98 Welsh cy Indo-European 1002 82.3 96.1
99 Wolof wo Atlantic-Congo 351 90.6 87.5
100 Xhosa xh Atlantic-Congo 1034 90.9 30.2
101 Yoruba yo Atlantic-Congo 816 92.4 84.6
102 Zulu zu Atlantic-Congo 822 85.5 67.2
All (102) 76.9 86.7
Table 6: Language name, code, family, and number of examples for each test set found in FLEURS. We report R@1 for mSLAM and PaLM 2 DE.